# Deanonymizing Ethereum Validators: The P2P Network Has a Privacy Issue

### v1

@@ 4.1 Ethereum Network-Layer Logging
We log all Ethereum network-layer messages by writing our own listening node implementation based on the Prysm client [57] (the most wide-spread consensus layer client for Ethereum [2]). We call our modified consensus node Rainbow throughput, as Rainbow acts as a prysm, taking peer input and breaking it down into the “colors” of the validators.
Rainbow connects to up to 1,000 peers, and statically subscribes to all subnets. These modifications allow us to deanoymize a larger set of nodes given that we have a higher peer count and observe attestations from all subnets. Other than that, it behaves as any Prysm node would, but logs three main sources of data. First and foremost, all attestations, their origin, and their origin subnet are logged. Second, this data is enriched by logging all advertised static subscriptions of our peers. Finally, we save precise connection data for all nodes we interact with.
Three instances of Rainbow are run alongside a Geth execution client on AWS r5a.4xlarge machines, in the us-east-1 (referred to as VA throughout), eu-central-1 (referred to as FR throughout), and ap-northeast-2 (referred to as SO throughout) data centers respectively. Additionally, we run one node on a bare-bones server in Zurich (referred to as ZH throughout). The data collection spans over three days from (May 7 00:00 UTC to May 10 00:00 UTC). About 700 GB of compressed data is collected and loaded into SQL databases for further processing.
Due to the gossiping nature of the network layer, the same attestations are often received from multiple sources. In this work, we consider only the attestation which we receive first.
@@ 4.2 Ethereum Network Coverage
To understand our coverage of the Ethereum Beacon network, we gather measurements of all nodes in the network via our own crawler implementation. Similarly to the Execution network of Ethereum (detailed in [43] and [33]), a node in the Beacon network keeps a peer table of nodes they have heard about, which they reference when they need to make more peer connections. Appendix B describes our crawler implementation which we use to enumerate through peer tables for all nodes we hear about in the network.
We run our crawler starting at 00:00 UTC on May 7, 8, and 9, and record all discovered ENRs***An ENR record encodes various information about a node including its network public key, IP, TCP/UDP ports, protocols supported and their versions, among others. for nodes running on the main Beacon network. We discover an average of 16.5K unique IPs (20K unique node IDs) running on the mainnet each day, with a cumulative 20,240 IPs (28,998 IDs) found in the three days of crawls.
Additionally, each hour we send an in-protocol ping message to all ENRs discovered in the previous crawl to distinguish nodes that are reachable in the network (i.e., are not behind a NAT or firewall and thus can accept incoming connections from our Rainbow nodes). We discover an average of 8.1K unique IPs (8.4K node IDs) that are reachable (at some point) each day, with a total of 8,941 IPs (9,468 IDs) of reachable nodes found in the three-day measurement period.
@@ 5.1 Peering Overview
Over the three-day data collection period, our four rainbow nodes attempted connections to 11,219 unique peers and successfully established connections with 10,785 (see Table 1). Finally, they held sufficiently long connections with 4,372 unique peers over the four rainbow nodes. Here, we consider a peer to have a long connection, if the total connection time over the three days lasted more than 32 epochs. Note that we disregard individual connections shorter than one epoch. Throughout, we consider a unique peer to be a unique IP port combination***Note that we used unique IPs in the crawler as not all ENRs provide a port, but for peer connections, we may connect to more than one node behind an IP but at different ports. Additionally, a node may change its public peer ID, though the same validators remain at the node..
@@
Lastly, we explore the unique peers we connected to for the period required for the analysis. We investigate the pairwise overlap across the Rainbow nodes in Table 2. We observe the biggest relative overlap between the FR and SO Rainbow nodes with 496 – around half of the peers these two nodes maintain long connections with overlap. Additionally, both the FR and SO nodes have an overlap of approximately 550 long peers with the VA node, while this constitutes more than half the long peers of the FR and SO nodes it only accounts for one-fourth of the VA long peers. Finally, the ZH node has the smallest overlap with the remaining nodes in relative and absolute terms. Possibly this is related to the three other nodes all being run from AWS and thus connecting to a more similar set of peers. Note that we provide further details on the connection lengths in Appendix C.
@@ 5.2 Deanonymization
No Validators. We did not receive a single non-backbone attestation from the peer and thus can safely assume that there are no validators hosted on the peer. ***This gives us a conservative lower bound on the number of validator nodes, and roughly matches previous analyses which estimate about half the network to be running at least one validator[13] - more details in Section 10
@@
Additionally, we note that the FR and SO nodes connect to a larger proportion of no-validator peers than the VA and ZH nodes. Thus, the VA and ZH nodes locate more validators on their peers, i.e., deanonymize more peers. We hypothesize that VA and ZH have higher-quality information and thus maintain more long-term peers and make it to more fanouts. This is a chicken-or-the-egg scenario, where being better connected leads to better connections given the greedy behavior of the network. It is unclear why VA and ZH have this advantage.
@@
Finally, we wish to highlight that out of the peers for which we do not conclude that they host no validators, our methods can deanonymize 84.42% across the four Rainbow nodes.
5.3 Validator Distribution
For those peers on which we deanonymize validators, we proceed by analyzing how many validators they host (see Figure 4). There is only a small number of validators located on most peers: a single validator on around 27% of our peers, with more than half of our peers having no more than 4 validators. In fact, we see that only 11% of peers have more than 100 validators. We further note that an in-detail inspection of Figure 4 reveals that for those peers that host many validators (i.e., more than 100), there is a bias towards a “round” number of validators to be hosted on a peer. This is evident by increases in the cdf at 100, 250, 300, 350, 400, 500, and 1,000 validators per peer. The gray vertical lines indicate these. For one, this indicates that large organizations running multiple validators on a machine tend to divide validators across nodes in round numbers as one would expect, e.g. an organization controlling 10,000 validators might divide them across 10 nodes and run 1,000 validators on each node. Additionally, this observation also indicates that our deanonymization technique locates all validators hosted on a machine as it is less likely for these “round” numbers to be a coincidence.
@@ 6 Verification
Overall, we locate 235,719 validators across the four Rainbow nodes – approximately a fourth of the validator set (see column one in Table 4). The ZH Rainbow node deanonymizes the majority with 191,164 while the VA node deanonymizes 78,708. Further, we deanonymize 15,276, and 14,199 validators with the FR and SO nodes respectively.
@@
Before we further analyze the insights into the Ethereum P2P network we can draw from our deanonymization, we first verify our deanonymization. To do so, we perform the following three steps: (1) check the consistency of validators hosted on the same machine, (2) analyze whether we uniquely identify validators to be hosted on a single machine, and (3) check whether we identify the same set of validators on a peer from multiple Rainbow nodes.
@@ 6.1 Consistency of Validator Sets
We start by investigating whether the set of validators our analysis concludes to be hosted on each peer is consistent with their attributes, e.g., they all belong to the same staking pool as opposed to belonging to five different staking pools. We consider a set of validators on a peer***Recall that a peer is a unique IP port combination. to be consistent, if one of the following holds:
@@
    We have entity labels for at least 30% of the validators and at least 90% of these are identical.***Note that we make five exceptions where we find validators from two staking pools that both enlist at least one common node operator on the same machine. In these cases, it appears that the node operator runs validators from different staking pools. We provide a discussion of this phenomenon in Section 7.2.
@@
The first condition is the most straightforward – we consider a validator set to be consistent if we have sufficiently many labels for the validator and they are largely identical. Note that we do not require them to be fully identical as there could be small inconsistencies in the entity label data set. Conditions 2 and 3 are very similar in that we require that the addresses used to deposit the stake and receive the execution layer rewards are largely identical for the validators in the set. Using the same address for either indicates that the validators belong to the same entity, i.e., the entity that controls the respective address. Finally, if the validators in the set have consecutive IDs this also indicates that they belong to the same entity, as the entity is likely to deposit the 32 ETH stake for a couple of validators in a row. This would be too much of a coincidence otherwise. We note that the peer set on all peers on which we only locate a single validator is trivially consistent.
@@
    We have entity labels for at least 10% of the validators, but less than 90% of these are identical (except if they are not all ENS names***Human-readable names associated with Ethereum addresses or Rocketpool***Rocketpool validators are often home run and an individual could run their personal and Rocketpool validators on the same machine.).
@@
Table 5 shows the consistency of the validator sets located on the peers, both by percent of each node’s peers and the proportion of located validators that are part of a consistent validator set. We highlight that the vast majority of validator sets located on peers are consistent – 93% of peers across all machines.
When we, however, turn to the proportion of validators these figures experience a significant shift – only 67.08% of validators located are part of consistent sets. This is a result of the inconsistent validator sets being significantly larger on average than the average validator set size. We proceed with a more in-depth investigation of these inconsistent validators.
@@
Figure 5: Pairwise overlap of the validator set of all peers with an inconsistent set of at least 20 validators. The darker the color of cell $(i,j)$, the larger the proportion of validators in the set of peer $j$ also found on peer $i$.
Service providers. There are a total of 15 peers on which we locate inconsistent validator sets containing at least twenty validators, with the biggest validator set having 47,595 validators. We calculate the pairwise overlap between these sets and visualize the result in Figure 5 with peers sorted by size (i.e., peer 0 has the largest validator set, and peer 14 has the smallest). Each peer has a non-zero overlap with at least one of the other peers, and there is significant overlap between many of the peers. While we cannot know for sure what the reason is, we presume that these peers are what we will refer to as P2P service providers throughout, i.e., service providers that help validators to quickly disseminate and receive messages e.g., bloXroute[10].
@@
Given that we suspect these 15 peers to be P2P service providers, i.e., they likely do not operate the validators but have priority access to their messages, we exclude them from our deanonymization. The second column in Table 4 indicates the number of validators we locate after excluding the aforementioned service providers. Overall we locate 154,591 validators – more than 15% of the network.
@@
6.2 Uniqueness of Validator-IP Mapping
To further verify our deanonymization we investigate the uniqueness of validator-IP port mappings – whether we only map a validator to a single IP or multiple ones. Column three in Table 4 indicates the number of validators with non-unique IP mapping. In total, there are 16,750 such non-unique validators, roughly 10% of all deanonymizations.
While this initially appears significant, a closer inspection suggests that we are observing some special behavior of peers in the network. Nearly three-fourths of this overlap is an almost full overlap between peers located in the same city. It appears that some entities, possibly as a means to ensure that their validators are always online, pass the message from their validators to multiple machines. As we will discuss in Section 9 such behavior can improve the privacy of validators. However, our deanonymization is still capable of tying these validators to the set of peers they run through that we are connected to. We have strong evidence to suggest such behavior for the majority of non-unique validator-IP mappings. The remaining overlap (around 2.5% of all deanonymized validators) is quite small and could be due to similar practices.
@@ 6.3 Similarity of Deanonymizations
The final analysis we perform is to verify our deanonymization is checking whether we locate the same set of validators on a peer (i.e., IP port combination) from multiple Rainbow nodes. In total, we maintained long connections to 791 peers from more than one Rainbow node. For 758 (95.83%) of these peers, we located the exact same set of validators on the machine from all Rainbow nodes that were connected to the peer. Further, on average the overlap of the validator sets is 99.09%. We conclude that our deanonymization appears robust – we locate the same validator set on a peer from different geographically distributed Rainbow nodes.

### v2

@@ 5.1 Ethereum Network-Layer Logging
We log all Ethereum network-layer messages by writing our own listening node implementation based on the Prysm client [66] (the most wide-spread consensus layer client for Ethereum [1]). We call our modified consensus node Rainbow throughout, as Rainbow acts as a Prysm, taking peer input and breaking it down into the “colors” of the validators.
Rainbow connects to up to 1,000 peers, and statically subscribes to all subnets. These modifications allow us to deanonymize a larger set of nodes given that we have a higher peer count and observe attestations from all subnets. Other than that, it behaves as any Prysm node would, but logs three main sources of data. First and foremost, all attestations, their origin, and their origin subnet are logged. Second, this data is enriched by logging all advertised static subscriptions of our peers. Finally, we save precise connection data for all nodes we interact with.
Three instances of Rainbow are run alongside a Geth execution client on AWS r5a.4xlarge machines, in the us-east-1 (referred to as VA throughout), eu-central-1 (referred to as FR throughout), and ap-northeast-2 (referred to as SO throughout) data centers respectively. Additionally, we run one node on a bare-bones server in Zurich (referred to as ZH throughout). The data collection spans three days from (May 7, 2024 00:00 UTC to May 10, 2024 00:00 UTC). About 700 GB of compressed data is collected and loaded into SQL databases for further processing.
Due to the gossiping nature of the network layer, the same attestations are often received from multiple sources. In this work, we only consider the attestation that we receive first.
@@ 5.2 Ethereum Network Coverage
To understand our coverage of the Ethereum Beacon network, we gather measurements of all nodes in the network via our own crawler implementation.
Similarly to the execution network of Ethereum (detailed in [47] and [35]), a node in the Beacon network keeps a peer table of nodes they have heard about, which they reference when they need to make more peer connections. The peer table is based on the Kademlia hash table [56] where each node stores ENRs [^9] An ENR record encodes various information about a node including its network public key, IP, TCP/UDP ports, protocols supported and their versions, among others. into buckets containing nodes whose ids (a 256-bit unique identifier computed from the ENR) has an XOR distance of $i$ from itself for $2560\leq i<$ 256. Nodes periodically update their peer tables using the discv5 peer discovery protocol [33]. We leverage the discovery protocol messages to enumerate a node’s peer table, and repeat this for all nodes we hear about in the network.
We run our crawler starting at 00:00 UTC on May 7, 8, and 9, and record all discovered ENRs for nodes running on the main Beacon network. We discover an average of 16.5K unique IPs (20K unique node IDs) running on the mainnet each day, with a cumulative 20,240 IPs (28,998 IDs) found in the three days of crawls.
Additionally, each hour we send an in-protocol ping message to all ENRs discovered in the previous crawl to distinguish nodes that are reachable in the network (i.e., are not behind a NAT or firewall and thus can accept incoming connections from our Rainbow nodes), we show this in Figure 3.101010We note that the in-protocol ping messages use the node’s ENR value, and some nodes do not respond if the ENR used is not up-to-date. Since the ENR changes if any of the values encoded change (e.g., a change in subnet subscriptions), ENRs change with some frequency. The gradual drop of responses over time we see in Figure 3 can be explained by both general network churn and ENRs learned in the previous crawl becoming outdated. We discover an average of 8.1K unique IPs (8.4K node IDs) that are reachable (at some point) each day, with a total of 8,941 IPs (9,468 IDs) of reachable nodes found in the three-day measurement period.
@@ 6.1 Peering Overview
Over the three-day data collection period, our four rainbow nodes attempted connections to 11,219 unique peers and successfully established connections with 10,785 (see Table 1). Finally, they held sufficiently long connections with 4,372 unique peers over the four rainbow nodes, representing approximately half of the reachable network (see Section 5.2) and one-third of the estimated total network size [43]. Here, we consider a peer to have a long connection, if the total connection time over the three days lasted more than 32 epochs. Note that we disregard individual connections shorter than one epoch. Throughout, we consider a unique peer to be a unique IP port combination. [^11] Note that we used unique IPs in the crawler as not all ENRs provide a port, but for peer connections, we may connect to more than one node behind an IP but at different ports. Additionally, a node may change its public peer ID, though the same validators remain at the node.
@@
Lastly, we explore the unique peers we connected to for the period required for the analysis. We investigate the pairwise overlap across the Rainbow nodes in Table 2. We observe the biggest relative overlap between the FR and SO Rainbow nodes with 496 – around half of the peers these two nodes maintain long connections with overlap. Additionally, both the FR and SO nodes have an overlap of approximately 550 long peers with the VA node, while this constitutes more than half the long peers of the FR and SO nodes it only accounts for one-fourth of the VA long peers. Finally, the ZH node has the smallest overlap with the remaining nodes in relative and absolute terms. Possibly this is related to the three other nodes all being run from AWS and thus connecting to a more similar set of peers.
@@ 6.2 Deanonymization
No Validators. We did not receive a single non-backbone attestation from the peer and thus can safely assume that there are no validators hosted on the peer. [^12] This gives us a conservative lower bound on the number of validator nodes, and roughly matches previous analyses which estimate about half the network to be running at least one validator[14] - more details in Section 10.
@@
Note that the FR and SO nodes connect to a larger proportion of no-validator peers than the VA and ZH nodes. As a result, the VA and ZH nodes can locate more validators on their peers, i.e., deanonymize more peers. We hypothesize that VA and ZH have higher-quality information and thus maintain more long-term peers and make it to more fanouts. This is a chicken-or-the-egg scenario, where being better connected leads to better connections given the greedy behavior of the network. It is unclear why VA and ZH have this advantage.
@@
Finally, we wish to highlight that out of the peers for which we do not conclude that they host no validators, our methods can deanonymize 84.57% across the four Rainbow nodes.
6.3 Deanonymizations Over Time and Across Rainbow Nodes
Next, we explore the impact of attack duration and the number of machines on the portion of the network that can be deanonymized.
In Figure 5, we look at the cumulative unique new peers our Rainbow nodes connected to with long connections over the measurement period. We observe an almost linear increase in the cumulative number of new long connections, two-fold the amount for the VA machine with most new connections. This highlights the advantage of deploying multiple nodes to increase the number of validators located in the network.
Figure 5: The cumulative number of peers each of our Rainbow nodes connected to over time, as well as the cumulative count across all.
Additionally, we also look at the number of validators deanonymized by our nodes over time in Figure 6. We perform our deanonymization on six hour fragments of our data, i.e., we analyze the first 6, 12, 18, and so forth, hours of data up until the entire collection period. Note that the number of deanonymized peers can decrease over time, which is particularly noticeable for the SO Rainbow node. This may occur if an initially strong connection to a peer weakens as time passes such that it no longer satisfies our heuristics. [^13] Such behavior might be observed if a peer removes us from a fanout. While a more careful analysis could consider only the best connection period for each peer, for a more cautious analysis we omit these from the numbers in the previous section.
In Figure 6, we observe a sharp initial increase in the number of deanonymized validators [^14] Note that this figure excludes validators located on P2P service provider. See Section 7.1 for details. during the first one and a half days. Although this trend slows significantly over time, the total number of deanonymized validators continues to grow throughout the measurement period. This growth is smaller than that of new peer connections (see Figure 5) as we come across more nodes that are not running validators.
Figure 6: The number of validators (excl. service providers) deanonymized by our Rainbow nodes over time, as well as the cumulative count across all. Note that the number of deanonymized peers can decrease over time as an initially good connection to a peer can degrade over time. This is particularly evident for the SO Rainbow node.
Overall, we conclude that various factors impact the success of deanonymization. A longer time horizon and deploying more nodes are generally advantageous, as highlighted in our analysis. Other factors, such as geographical location and hardware configurations (e.g., bare-metal servers versus virtual machines), significantly influence a node’s performance within the network [47]. Low-latency connections in particular play a critical role in determining the performance of a node in the network. At the same time, networks are inherently complex, and a node’s performance can vary unpredictably.
An attacker aiming to deanonymize the most validators would benefit from deploying multiple nodes in the network over an extended time horizon. The associated costs depend on how the nodes are deployed. Our measurements were relatively expensive, as we operated three nodes using AWS machines, with network traffic being the primary expense. [^15] Each node incurred a cost of approximately 100 US\$ per day. However, in general, a residential internet connection is sufficient, and the hardware requirements for Ethereum full nodes are modest [4] (less than 1,000 US\$). While these expenses are not trivial, they are minor compared to the potential profits from such attacks. For context, the total MEV extracted since the Ethereum Merge in September 2021 amounts to 598,019 ETH (approximately 2 billion US\$) [78].
Furthermore, an attacker could optimize costs by refining their strategy – for example, rotating through peers instead of keeping long-term connections to peers on which validators are already deanonymized.
@@ 7 Verification
Overall, we locate 252,293 validators across the four Rainbow nodes – approximately a fourth of the validator set (see column one in Table 4). The ZH Rainbow node deanonymizes the majority with 215,293 while the VA node deanonymizes 74,904. Further, we deanonymize 13,771, and 14,388 validators with the FR and SO nodes respectively.
@@
Before delving deeper into the insights about the Ethereum P2P network derived from our deanonymization, we first validate the deanonymization itself. Notably, while we lack access to ground truth data, we employ several methods to assess the accuracy of our results. Although these methods are not definitive, they collectively provide valuable insights into the quality of our deanonymization. Specifically, we undertake the following three steps: (1) verify the consistency of validators hosted on the same machine, (2) analyze whether validators can be uniquely identified as being hosted on a single machine, and (3) confirm whether the same set of validators is identified on a peer across multiple Rainbow nodes.
@@ 7.1 Consistency of Validator Sets
We start by investigating whether the set of validators our analysis concludes to be hosted on each peer is consistent with their attributes, e.g., they all belong to the same staking pool as opposed to belonging to five different staking pools. We consider a set of validators on a peer [^16] Recall that a peer is a unique IP port combination. to be consistent, if one of the following holds:
@@
    We have entity labels for at least 30% of the validators and at least 90% of these are identical. [^17] Note that we make five exceptions where we find validators from two staking pools that both enlist at least one common node operator on the same machine. It appears that the node operator runs validators from different staking pools. We provide a discussion of this phenomenon in Section 8.3.
@@
The first condition is the most straightforward – we consider a validator set to be consistent if we have sufficiently many labels for the validator and they are largely identical. Note that we do not require them to be fully identical as there could be small inconsistencies in the entity label data set. Conditions 2 and 3 are very similar in that we require that the addresses used to deposit the stake and receive the execution layer rewards are largely identical for the validators in the set. Using the same address for either indicates that the validators belong to the same entity, i.e., the entity that controls the respective address. Finally, if the validators in the set have consecutive IDs this also indicates that they belong to the same entity, as the entity is likely to deposit the 32 ETH stake for a couple of validators in a row. This would otherwise be too much of a coincidence. Finally, a peer where we locate only a single validator is trivially consistent.
@@
    We have entity labels for at least 10% of the validators, but less than 90% of these are identical (except if they are not all ENS names [^18] Human-readable names associated with Ethereum addresses or Rocketpool [^19] Rocketpool validators are often home run and an individual could run their personal and Rocketpool validators on the same machine.).
@@
Table 5 shows the consistency of the validator sets located on the peers, both by percent of each node’s peers and the proportion of located validators that are part of a consistent validator set. We highlight that the vast majority of validator sets located on peers are consistent – 93.75% of peers across all machines.
When we, however, turn to the proportion of validators these figures experience a significant shift – only 63.67% of validators located are part of consistent sets. This is a result of the inconsistent validator sets being significantly larger on average than the average validator set size. We proceed with a more in-depth investigation of these inconsistent validators.
@@
Figure 7: Pairwise overlap of the validator set of all peers with an inconsistent set of at least 200 validators. The darker the color of cell $(i,j)$, the larger the proportion of validators in the set of peer $j$ also found on peer $i$.
Service providers. There are a total of 17 peers on which we locate inconsistent validator sets containing at least 200 validators, with the biggest validator set having 84,165 validators. We calculate the pairwise overlap between these sets and visualize the result in Figure 7 with peers sorted by size (i.e., peer 0 has the largest validator set, and peer 16 has the smallest). Each peer has a non-zero overlap with at least one of the other peers, and there is significant overlap between many of the peers. We presume that these peers are what we will refer to as P2P service providers throughout, i.e., service providers that help validators to quickly disseminate and receive messages e.g., bloXroute[11].
@@
Given that we suspect these 17 peers to be P2P service providers, i.e., they likely do not operate the validators but have priority access to their messages, we exclude them from our deanonymization. The second column in Table 4 indicates the number of validators we locate after excluding the aforementioned service providers. Overall we 161,057 validators – more than 15% of the network.
7.2 Uniqueness of Validator-IP Mapping
@@
To further verify our deanonymization we investigate the uniqueness of validator-IP port mappings – whether we only map a validator to a single IP or multiple ones. Column three in Table 4 indicates the number of validators with non-unique IP mapping. In total, there are 16,172 such non-unique validators, roughly 10% of all deanonymizations.
While this initially seems significant, closer inspection reveals special behavior by certain peers. Nearly three-fourths of the overlap involves peers in the same city, suggesting that some entities, possibly to ensure validator uptime, relay messages from their validators through multiple machines. As discussed in Section 9, such practices can enhance validator privacy. However, our deanonymization still ties these validators to the peer sets we connect to. We find strong evidence of this behavior for many non-unique validator-IP mappings. The remaining overlap, less than 1% of all deanonymized validators (excluding service providers), is minimal and may result from similar practices.
@@ 7.3 Similarity of Deanonymizations
The final analysis we perform is to verify our deanonymization is checking whether we locate the same set of validators on a peer (i.e., IP port combination) from multiple Rainbow nodes. In total, we deanonymized validators from more than one Rainbow node for 794 peers. For 762 (95.96%) of these peers, we located the exact same set of validators on the machine from all Rainbow nodes that were connected to the peer. Further, on average the overlap of the validator sets is 99.20%. Thus, our deanonymization appears robust – we locate the same validator set on a peer from different geographically distributed Rainbow nodes.